# SoliDiffy: AST Differencing for Solidity Smart Contracts

### v3

@@ V-B Results for RQ2
Figure 6 shows the results for the two differencing tools while running the diff pairs with the same operators repeatedly applied on the same Solidity contract (outlined in Section IV-C). The results are presented as a violin plot at which the width of the violin at different points shows the density of data. Key statistical markers, such as the mean and median in this plot highlight the skewness of the data. Peaks in the violin indicate where data clusters, and the tails represent outliers or extreme values. As Figure 6 shows, SoliDiffy is more dense towards the lower values of edit scripts, especially for lower number of mutations used for contract pair generation. For instance, differencing task performed on contract pairs which were the result of one and two mutations are very dense towards very low values of edit scripts for SoliDiffy.
@@
Result for RQ2: SoliDiffy consistently produces smaller edit distances compared to Difftastic, regardless of the number of modifications in the smart contracts to be diffed. The finding holds statistical significance.
V-C Results for RQ3
@@
We conducted a Wilcoxon signed-rank test [39] to assess the statistical significance of the observed differences between SoliDiffy and Difftastic across the $44$ mutation operators. The test results show that for all $44$ operators, the differences between SoliDiffy and Difftastic were statistically significantThis confirms that the visual performance differences seen in Figure 7 are not due to random chance.
@@ V-D Results for RQ4
Result for RQ4: SoliDiffy works well to analyze open-source commits changing real-world Solidity smart contracts. SoliDiffy is readily usable by researchers and practitioners for advanced smart contract analysis tasks.
VI Discussion
VI-A Lessons Learned
@@
VI-B Threats to Validity
To evaluate SoliDiffy , we used mutants for generating diff pairs. This approach provided fine-grained control over the differences in the code, enabling us to systematically evaluate the tool (see Sections V-A, V-B, V-C). By using mutants instead of relying on the commit history of Solidity project repositories, we ensured that each diff pair was isolated and clearly attributable to specific types of syntactic changes. This level of control facilitated a more detailed analysis of the tools’ behavior under specific conditions.
However, using mutants also introduces some potential problems. Firstly, the differences between the resulting files may be identical across some casesFurthermore, many mutation operators generated minor changes, which, when analyzed individually, provided limited insight into the broader performance of the differencing tools. This could lead to an underestimation of the tools’ effectiveness in handling more complex, real-world code changes.
Even in cases where the large-scale evaluation showed differences in edit distances between the tools, these differences were not always apparent in the visual representation of the diffs.
Statistical analysis on edit distances also may introduce potential threats to validity. First, while non-parametric tests like the Wilcoxon signed-rank and Kolmogorov-Smirnov (K-S) tests are suitable for our non-normal data distributions, they assume comparable scales across tools. Differences in edit script lengths between SoliDiffy and Difftastic may partly reflect varying granularities rather than inherent tool efficacy. Second, the dataset’s synthetic nature (using controlled mutants) could bias statistical comparisons, as the tool behaviors observed in a laboratory setting may not fully extend to real-world contract complexities. Finally, the choice of statistical significance alone may overlook practical differences. While the tests confirm tool distinctions, additional metrics such as effect size or real-world user studies could provide further insights into the impact of these differences on developer tasks.
VII Related Work
@@
IJM [45] Merges nodes, prunes sub-trees for faster differencing Improved GumTree algorithm Shorter, faster edit scripts
@@
Yang & Whitehead’s Pruning [47] Pruning ASTs based on unchanged lines Line-based pruning for AST Pruning has negligible impact on diff results
@@
Difftastic [21] Supports Solidity Text changes differencing Lacks concrete evaluation, limited to text-based diffs, unsuitable for complex diff tasks, Supports only two edit actions
@@
SoliDiffy (current work) Move operation in edit scripts, RTED algorithm, Solidity-specific differencing Top-down and bottom-up AST traversal and mapping to generate edit script for a diff pair, supports four edit actions Issues with visual representation of diffs, Edit script length for very small changes
As Solidity source code differencing is a contributing technical piece of many solutions such as version control [48], automated debugging [5], and code repair [49, 50], we provide a detailed review of tools and research on source code differencing.
VII-A Code Differencing
Tools such as srcML [25] and Dex [41] proposed ways to enhance code differencing, while UMLDiff [42] and Diff/TS [43] aimed at improving accuracy and efficiency. Meanwhile, OperV [20] explored varying levels of granularity in differencing.
The srcML tool converts source code into an XML-based intermediate representation, retaining both the syntax and textual elements, allowing for regenerating the original code [25]. Unlike AST-based tools, it uses the Unix diff command on these XML files.
Dex introduced the use of Abstract Semantic Graphs (ASGs) instead of ASTs for C code, linking related nodes like variable references and declarations [41]. Its differencing approach used graph rather than tree differencing, excluding the move operation from its edit scripts, but achieving a 95% accuracy in detecting correct edit actions.
UMLDiff [42] and Diff/TS [43] focused on enhancing the structural analysis of code changes. UMLDiff used class models reverse-engineered from Java source code to build change trees and calculated similarity scores to detect changes in the overarching class structure [42]. Diff/TS, on the other hand, combined tree differencing with configurable heuristics to improve runtime and accuracy, incorporating all standard edit actions in its scripts [43].
OperV sought to offer variable granularity in version control systems, combining line-based differencing with AST-based matching, though its evaluation was limited compared to more modern approaches [20].
@@ VII-A Code Differencing
Recent AST differencing tools, including GumTree [19], MTDIFF [44], IJM [45], and the approach by Matsumoto et al. [46], focus on refining the generation of edit scripts (Section II-B2).
One common method for evaluating edit script quality is by measuring its length [19, 46, 44]. Shorter edit scripts are generally preferred because they tend to contain fewer redundant operations and more closely align with the actual code modifications. Another approach is to count the number of matched nodes in the initial differencing step, which are not included in the edit script, providing insight into the tool’s effectiveness in detecting unchanged code structures. However, there are criticisms of these methods. For instance, focusing solely on reducing script length can sometimes lead to suboptimal results, as seen in tools like SrcDiff [22].
@@
VII-B Gumtree Family of AST Differencing.
The Changedistiller algorithm [52] is a foundational work in AST differencing, introducing a method to match identical nodes between two ASTs and generate an edit script. It built upon Chawathe’s 1996 algorithm [28], optimizing it for source code by reducing edit script length by 45%. This approach influenced many subsequent tools, including GumTree [19].
GumTree is particularly notable for its introduction of the move operation in edit scripts, which improves accuracy by grouping related changes. It uses a combination of top-down and bottom-up AST traversal and incorporates the RTED algorithm [53] for generating mappings in smaller sub-trees. However, GumTree has not integrated the more recent APTED algorithm [54], which could enhance its performance.
GumTree also supports hyperparameter tuning to optimize edit script length, as demonstrated by Martinez’s Diff Auto Tuning (DAT) technique, which reduced script length [55]. Additionally, GumTree can process general-purpose Tree-sitter CSTs by converting them into a format suitable for AST differencing [56].
MTDIFF [44] and Iterative Java Matcher (IJM) [45] introduced improvements on built on top of core GumTree algorithm but lack maturity and integration into the mainstream differencing code tool.
While the aforementioned tools offer improvements for general-purpose languages, they lack specific adaptations for Solidity. SoliDiffy builds on this line of work by building on top of Gumtree’s algorithms and ecosystem for Solidity smart contracts.
VII-C Solidity Code Differencing
Research on code differencing specific to Solidity is limited. While line-based differencing can be used across languages, it lacks the precision needed for Solidity’s unique syntax and semantics, as discussed in Section II-B.
The only dedicated Solidity differencing tool in the literature is part of the Solidity Instrumentation Framework (SIF) [57]. SIF uses AST-differencing, but its implementation is poorly documented and relies on an outdated AST format no longer supported by the Solidity compiler, making it unusable for newer code.
Outside academic literature, Smartdiffer [58] links Etherscan with a line-based differencing tool but does not perform AST differencing, limiting its relevance to more sophisticated analysis.
Difftastic [21] is a more complete AST differencing tool that supports Solidity. It uses a tree-sitter parser to generate side-by-side diffs or JSON output. However, it does not use traditional edit script generation, instead focusing on concrete text changes, and lacks thorough research-based evaluation.
VIII Conclusion

### v4

@@ VI-B Results for RQ2
Figure 6 shows the results for the two differencing tools while running the diff pairs with the same operators repeatedly applied on the same Solidity contract (outlined in Section V-C). The results are presented as a violin plot at which the width of the violin at different points shows the density of data. Key statistical markers, such as the mean and median in this plot highlight the skewness of the data. Peaks in the violin indicate where data clusters, and the tails represent outliers or extreme values. As Figure 6 shows, SoliDiffy is more dense towards the lower values of edit scripts, especially for lower number of mutations used for contract pair generation. For instance, differencing task performed on contract pairs which were the result of one and two mutations are very dense towards very low values of edit scripts for SoliDiffy.
@@
Result for RQ2: SoliDiffy consistently produces smaller edit distances compared to Difftastic, regardless of the number of modifications in the smart contracts.
VI-C Results for RQ3
@@
We conducted a Wilcoxon signed-rank test [36] to assess the statistical significance of the observed differences between SoliDiffy and Difftastic across the $44$ mutation operators. The test results show that for all $44$ operators, the differences between SoliDiffy and Difftastic were statistically significant. This confirms that the visualized differences seen in Figure 7 are not due to random chance.
@@ VI-D Results for RQ4
Result for RQ4: SoliDiffy provides structured and accurate differencing on real-world Solidity smart contract commits, significantly outperforming Git line differencing. The statistical tests confirm its reliability for practical use in tracking contract evolution.
VII Discussion
VII-A Lessons Learned
@@
IJM [44] Merges nodes, prunes sub-trees for faster differencing Improved GumTree algorithm Lacks integration with mainstream Gumtree
@@
Hunk-based AST Pruning [46] Pruning ASTs based on unchanged lines Line-based pruning for AST Pruning has negligible impact on diff results
@@
Difftastic [20] Supports Solidity Text changes differencing Lacks concrete evaluation; uses text-based diffs; and only supports two edit actions
@@
SoliDiffy Move operation in edit scripts, RTED algorithm, Solidity-specific differencing Top-down and bottom-up AST traversal and mapping to generate edit script for a diff pair, supports four edit actions Issues with visual representation of diffs, edit script length for very small changes
VII-B Threats to Validity
Construct Validity.
We evaluated SoliDiffy primarily through edit script length, a widely used metric in AST differencing research [16, 45]. While shorter edit scripts generally indicate more precise differencing, this metric does not capture all aspects of quality, such as human interpretability or the practical utility of the generated edit scripts for downstream tasks.
Internal Validity.
Our large-scale evaluation was performed on a synthetic dataset derived from the DAppScan [28] project. The use of mutation-based transformations ensured systematic variation but may not fully capture the complexity of organic contract evolution. To mitigate this, we supplemented our evaluation with real-world commits of Uniswap, but further studies across additional Solidity projects would provide deeper insights.
External Validity.
The effectiveness of SoliDiffy was demonstrated across a diverse range of contract pairs, including both synthetic and real-world data. However, the generalizability of our results may be influenced by the dataset’s composition. While Solidity is the dominant smart contract language, its ecosystem is rapidly evolving, and new language features or developer practices may introduce unforeseen challenges for AST differencing tools. Our open-source implementation allows the research community to extend our approach to newer Solidity versions and potentially adapt it for other smart contract platforms.
Conclusion Validity.
The statistical tests performed in our study confirm the significance of SoliDiffy’s improvements over existing tools. However, statistical significance does not always translate directly to practical performance improvement. While SoliDiffy consistently produced shorter edit scripts and handled more contract pairs successfully, the impact of these improvements on downstream tasks, such as security auditing and automated program repair, requires further investigation.
VIII Related Works
Here we provide a detailed review of the tools and research on source code differencing.
VIII-A Code Differencing From a Historical Perspective
The srcML [38] converts source code into an XML-based intermediate representation, retaining both the syntax and textual elements which allows for regenerating the original code. Unlike AST-based tools, it uses the Unix diff command on these XML files. Dex introduced the use of Abstract Semantic Graphs (ASGs) instead of ASTs for C code, linking related nodes like variable references and declarations [39]. Its differencing approach used graph rather than tree differencing, excluding the move operation from its edit scripts, but achieving a 95% accuracy in detecting correct edit actions. UMLDiff [40] and Diff/TS [41] focused on structural analysis of code changes. UMLDiff used class models reverse-engineered from Java source code to build change trees and calculated similarity scores to detect changes in the overarching class structure [40]. Diff/TS, on the other hand, combined tree differencing with configurable heuristics to improve runtime and accuracy, incorporating all standard edit actions in its scripts [41]. OperV sought to offer variable granularity in version control systems, combining line-based differencing with AST-based matching, though its evaluation was limited compared to more modern approaches [42].
@@ VIII-A Code Differencing From a Historical Perspective
Recent AST differencing tools, including GumTree [16], MTDIFF [43], IJM [44], and the approach by Matsumoto et al. [45], focus on refining the generation of edit scripts (Section III-B2). One common method for evaluating edit script quality is by measuring its length [16, 45, 43]. Shorter edit scripts are generally preferred because they tend to contain fewer redundant operations and more closely align with the actual code modifications. Another approach is to count the number of matched nodes in the initial differencing step, which are not included in the edit script, providing insight into the tool’s effectiveness in detecting unchanged code structures. However, there are criticisms of these methods. For instance, focusing solely on reducing script length can sometimes lead to suboptimal results, as seen in tools like SrcDiff [49].
@@
VIII-B Gumtree Family of AST Differencing.
The Changedistiller algorithm [51] is a foundational work in AST differencing, introducing a method to match identical nodes between two ASTs and generate an edit script. It built upon Chawathe’s 1996 algorithm [26], optimizing it for source code by reducing edit script length by 45%. This approach influenced many subsequent tools, including GumTree [16]. GumTree is particularly notable for its introduction of the move operation in edit scripts, which improves accuracy by grouping related changes. It uses a combination of top-down and bottom-up AST traversal and incorporates the RTED algorithm [52] for generating mappings in smaller sub-trees. GumTree also supports hyperparameter tuning to optimize edit script length, as demonstrated by Martinez’s Diff Auto Tuning (DAT) technique, which reduced script length [53]. Additionally, GumTree can process general-purpose Tree-sitter CSTs by converting them into a format suitable for AST differencing [54]. MTDIFF [43] and Iterative Java Matcher (IJM) [44] introduced improvements on built on top of GumTree algorithm but lack maturity and integration into the mainstream differencing tool. While the aforementioned tools offer improvements for general-purpose languages, they lack specific adaptations for Solidity. SoliDiffy builds on this line of work by building on top of Gumtree’s algorithms and ecosystem for Solidity smart contracts.
VIII-C Solidity Code Differencing
Research on code differencing specific to Solidity is limited. While line-based differencing can be used across languages, it lacks the precision needed for Solidity’s unique syntax. The only dedicated Solidity differencing tool in the literature is part of the Solidity Instrumentation Framework (SIF) [55]. SIF uses AST-differencing, but its implementation is poorly documented and relies on an outdated AST format no longer supported by the Solidity compiler, making it unusable for newer code. Outside academic literature, Difftastic [20] is the only other usable differencing tool that supports Solidity. It uses a Tree-sitter [27] parser to generate side-by-side diffs or JSON output. However, it does not use traditional edit script generation, instead it focuses on concrete text changes (word-by-word), and as of now, no thorough evaluation of Difftastic has been published..
IX Conclusion